My math problem says
Jessica goes to a store to buy $70 worth of flour and sugar for her bakery. A bag of flour costs $5, and a bag of sugar costs $7.
a. Write an equation to find the number of bags of each type Jessica can buy.
b. Graph your equation.
Thank you so much. Can you please show your work thank you!

Answer by rwm(914)

The hen and buffalo problem is very similar but had more info.
5f+7s=70
we don't know the total number of bags bought.
We have two unknowns and one equation. Not enough info
but notice that
5*7=35
7*5 =35
35+35=70
we could have 7 $5 bags and 5 $7dollar bags